# Discrete Chi-Square Method beats Discrete Fourier Transform and other similar parametric time series analysis methods

### v3

@@ 3.8 Best model
We conclude that the best model for the real data can be found by applying the Fisher-test to any arbitrary number of different nested DCM or DFT models.
3.9 Significance estimates
The Fisher-test gives the critical level $Q_{\mathrm{F}}$ for rejecting the $H_{0}$ hypothesis (Equation 40). After the detection of the first strongest signal, the values $Q_{\mathrm{F}}$ of these critical levels increase for the next detected weaker signals. Typical examples can be found in Jetsu 2025 (Tables S5-S16). No new signals are detected when $Q_{\mathrm{F}}>\gamma=0.001$ because the critical level exceeds the pre-assigned significance level and the $H_{0}$ hypothesis is no longer rejected.
The Fisher test critical level $Q_{\mathrm{F}}$ is the significance estimate for each signal detected after the first signal. However, this Fisher-test process gives no significance estimate for the first detected signal.
We estimate the critical level $Q_{\mathrm{F}}$ for the first detected signal $h_{1}(t)$ by testing the alternative that there are no signals in the data. The “no signal hypothesis” is
$h(t)=0\Rightarrow h_{1}(t)=0\Rightarrow g(t)=p(t).$ (41)
In Model 1 simulated data combination $n=100$ and ${\mathrm{SN}}=100$, the correct model is $g_{1,1,0}(t)$. We use the Fisher-test to compare this correct $g_{1,1,0}(t)$ model to different polynomial models $g(t)=p(t)$ having $K_{3}=0,1,..,7$ (Table 9).
The $\mathcal{M}$=1, 2 and 3 models have less free parameters than the correct $g_{1,1,0}(t)$ model, but the $\chi^{2}$ values of these three polynomial models are so large that the Fisher-test is merely a formality. The $\mathcal{M}$=4 model has the same number of free parameters as the correct $g_{1,1,0}(t)$ model, but the comparison of $\chi^{2}$ values reveals that this third order $p(t)$ polynomial is not the correct model for the data. The next model $\mathcal{M}$=5 has more free parameters than the correct $g_{1,1,0}(t)$ model. This fourth order $p(t)$ polynomial $\mathcal{M}$=5 model must be rejected because it has a larger $\chi^{2}$ value than the correct $g_{1,1,0}(t)$ model. The critical levels $Q_{\mathrm{F}}\gg\gamma=0.001$ for the remaining $\mathcal{M}$=6, 7 and 8 polynomial models are so large that these models must also be rejected. The results in Table 9 confirm the “no signal hypothesis” (Equation 41) rejection. The analysed Model 1 simulated data must contain at least the $h_{1}(t)$ pure sine signal. For the constant, linear or parabolic $p(t)$ trend alternatives, the critical level for this $h_{1}(t)$ signal is $Q_{\mathrm{F}}<\!10^{-16}$.
There is no need to discuss the DFT significance estimates [19, their Equation 22] because this method fails to detect the correct Model 1 period.
Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for predictive data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations). (1) Model. (2) Predictive data test statistic $z$ (Equation 16). (3) Predicted data test statistic $z_{\mathrm{pred}}$ (Equation 3.10). (4) Data file. (5) Control file. (6) Figure where prediction is shown.
@@ 3.9 Significance estimates
Model $z$ $z_{\mathrm{pred}}$ Data file Control file Figure
@@
Figure 8: DCM model prediction (Model 3 simulated data for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are predictive data). (a) One signal model $g_{1,1,0}$ results. Predictive data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and predicted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots). Continuous black line of predictive data model $g(t,\bm{\beta}_{\mathrm{pred}})$ ends to vertical dotted blue line, where dotted black line of predicted model $g^{\prime}(t,\bm{\beta}_{\mathrm{pred}})$ begins. Dotted red line denotes $\pm 3\sigma_{g(t)}$ errors of both models (Equation 22). Residuals of predictive data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of predicted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line). (b) Two signal model $g_{2,1,0}$ results. Otherwise as in “a”. (c) Three signal model $g_{3,1,0}$ results. Otherwise as in “a”.
3.10 Prediction
@@
There are numerous techniques for predicting (forecasting) a time series [e.g., 18, 20, 23]. The DCM model $g(t_{i},\bm{\beta})$ can be used to predict. We divide all data into the predictive data and the predicted data. The time points, the observations and the errors of these samples are
  • $n$ predictive data values $t_{i}$, $y_{i}$ and $\sigma_{i}$
  • $n^{\prime}$ predicted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$
The DCM gives the best predictive data model
$g_{i}=g(t_{i},\bm{\beta}_{\mathrm{pred}}),$ (42)
@@
where $\bm{\beta}_{\mathrm{pred}}$ are the free parameter values. The $t_{\mathrm{mid,pred}}\!=\!(t_{n}\!+\!t_{1})/2$ and $\Delta T_{\mathrm{pred}}\!=\!t_{n}\!-\!t_{1}$ values are computed from the predictive data time points $t_{i}$.
The predicted data model values are
$g^{\prime}_{i}$ $=$ $g(t^{\prime}_{i},\bm{\beta}_{\mathrm{pred}}),$ (43)
@@
where $t_{\mathrm{mid}}=t_{\mathrm{mid,pred}}$ and $\Delta T=\Delta T_{\mathrm{pred}}$. We do not compute “new” $t_{\mathrm{mid}}$ and $\Delta T$ values from the predicted data time points $t^{\prime}_{i}$ because the correct $g^{\prime}_{i}$ values are obtained only from the $\bm{\beta}_{\mathrm{pred}}$, $t_{\mathrm{mid,pred}}$ and $\Delta T_{\mathrm{pred}}$ combination of Equation 42. The $n^{\prime}$ predicted data model residuals
@@
give the predicted data test statistic
$z_{\mathrm{pred}}$ $=$ z test statistic for predicted
data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ (Equation 15 or 16).
@@
This parameter $z_{\mathrm{pred}}$ measures how well the prediction (Equation 43) obtained from the predictive data works for the predicted data. If the DCM detects a new signal from the predictive data, there are two alternatives:
@@
  • If this new signal is real, the $z_{\mathrm{pred}}$ value of predicted data decreases.
@@
  • If this new signal is unreal, the $z_{\mathrm{pred}}$ value of predicted data increases.
@@
This “Prediction-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
@@
We compute the $z_{\mathrm{pred}}$ parameter value from the known predicted data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ values (Equation 3.10) . Predictions are possible even if all $t^{\prime}_{i}$, $y^{\prime}_{i}$ or $\sigma^{\prime}_{i}$ predicted data values are unknown. In this case, the $t_{i}$, $y_{i}$ and $\sigma_{i}$ values of all data can be used as predictive data. The best DCM model $g(t_{i},\bm{\beta}_{\mathrm{pred}})$ for all data determines the correct $\bm{\beta}_{\mathrm{pred}}$, $t_{\mathrm{mid,pred}}=(t_{n}+t_{1})/2$ and $\Delta T_{\mathrm{pred}}=t_{n}-t_{1}$ combination. The $n^{\prime}$ predicted data time points $t_{i}^{\prime}$ can be created for any arbitrary chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{i}^{\prime}$. The $n^{\prime}$ predicted model $g_{i}^{\prime}=g^{\prime}_{i}(t_{i}^{\prime},\bm{\beta}_{\mathrm{pred}})$ values are obtained from Equation 43. These $g^{\prime}_{i}$ values can be used, for example, to compute the predicted mean level
$m_{\mathrm{pred}}={{1}\over{n^{\prime}}}\sum_{i=1}^{n^{\prime}}g^{\prime}_{i}$ (45)
@@
during the chosen time interval $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu (2025) used this $m_{\mathrm{pred}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720. Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@
  • If DCM detects a real new signal in all data, the $m_{\mathrm{pred}}$ value decreases.
@@
  • If DCM detects many real signals in all data, the $m_{\mathrm{pred}}$ value falls below $m$.
@@
  • If DCM detects an unreal new signal in all data, the $m_{\mathrm{pred}}$ value increases.
@@
We use Model 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data (Table 3, column 2) to illustrate the DCM prediction technique. The black dots in Figures 8a-c are the first $n=40$ predictive data values $t_{i}$, $y_{i}$ and $\sigma_{i}$. The open black dots denote the $n^{\prime}=10$ predicted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$. The continuous black line is the predictive model $g(t)$ and the dotted black line is the predicted model $g^{\prime}(t)$. The red dotted line shows the $\pm 3\sigma$ errors of both models. The blue dots are the predictive data $\epsilon_{i}=y_{i}-g_{i}$ residuals and the open blue dots are the predicted data $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ residuals.
@@
Model 3 is the sum of two $P_{1}=0.16$ and $P_{2}=0.17$ pure sine signals $(K_{1}=2,K_{2}=1)$ superimposed on the constant mean level $M_{0}=1$ $(K_{3}=0)$. We compute the $z$ and $z_{\mathrm{pred}}$ values for the $g_{1,1,0}$, $g_{2,1,0}$ and $g_{3,1,0}$ models (Table 10). These three models have the same $K_{2}=1$ and $K_{3}=0$ orders as Model 3, but their signal numbers $K_{1}=1,2$ or 3 are different, the $g_{2,1,0}$ model being the correct simulation Model 3.
@@
The correct $g_{2,1,0}$ model gives the smallest $z_{\mathrm{pred}}=0.88$ value (Table 10). Therefore, it is a better predictive model than the $g_{1,1,0}$ and $g_{3,1,0}$ models. The one signal $g_{1,1,0}$ model prediction fails because the blue open circles denoting the predicted data residuals $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ show large deviations from to the blue dotted line offset level of $\epsilon_{i}^{\prime}=-1$ (Figure 8a). The two signal $g_{2,1,0}$ model prediction succeeds because all open blue dots denoting the predicted data residuals stay close to the blue dotted line offset level $\epsilon_{i}^{\prime}=-1$, as well as inside the red dotted $\pm 3\sigma$ model error limits (Figure 8b).
@@
The three signal model $g_{3,1,0}$ periods are $P_{1}=0.0643\pm 0.0018$, $P_{2}=0.1592\pm 0.0012$ and $P_{3}=0.1716\pm 0.0012$. The amplitudes of these pure sine signals are $A_{1}=0.092\pm 0.025$, $A_{2}=1.879\pm 0.16$ and $A_{3}=1.82\pm 0.15$. The periods and the amplitudes of the two strongest $P_{2}$ and $P_{3}$ signals are correct because they are the same as in Model 3 simulation (Table 3, column 1). Therefore, the $g_{3,1,0}$ prediction appears nearly as good as the $g_{2,1,0}$ prediction because the residuals $\epsilon_{i}^{\prime}$ are close to the offset level of $\epsilon_{i}^{\prime}=-1$ (blue open dots), and these residuals also stay inside the red dotted $\pm 3\sigma$ model error limits (Figure 8c). The amplitude $A_{1}=0.092$ of the third $P_{1}=0.0643$ signal is very low. Due this weak “unreal” $P_{1}$ signal, the $g_{3,1,0}$ model has a larger $z_{\mathrm{pred}}=0.91$ value than the correct $g_{2,1,0}$ model (Table 10).
The $n=40$ predictive data parameters $\eta_{1}=7$, $\eta_{2}=10$, $\chi_{1}=22.946$ and $\chi_{2}=19.790$ for the simple $g_{2,1,0}$ model and the complex $g_{3,1,0}$ model give Fisher-test values $F_{\chi}=1.32$ and $Q_{\mathrm{F}}=0.28>\gamma_{\mathrm{F}}=0.001$. The $H_{0}$ hypothesis is not rejected. For the predictive data, the Fisher-test also confirms that the $g_{2,1,0}$ model is better than the $g_{3,1,0}$ model. Finally, we note that the three signal model $g_{3,1,0}$ is not unstable ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”) although the simulated data contains only two signals.
For the $g_{2,1,0}$ model, the DCM detects the correct Model 3 signal frequencies (Table 3), but these frequencies are “too close” for the correct DFT detection (Equation 29). Therefore, the DFT prediction can not succeed.
@@
The analyses of all seven complex data sets indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7). This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [22, 15] would explain why correct signal periods have not been detected earlier from the sunspot data.

### v4

@@ 3.10 Significance estimates
Table 10: DCM time series analysis between $P_{\mathrm{min}}=0.53$ and $P_{\mathrm{max}}=4.80$ for forecasting data (Model = 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data: first forty observations). (1) Model. (2) Forecasting data test statistic $z$ (Equation 17). (3) Forecasted data test statistic $z_{\mathrm{Fore}}$ (Equation 3.12). (4) Data file. (5) Control file. (6) Figure where forecast is shown.
@@
Model $z$ $z_{\mathrm{Fore}}$ Data file Control file Figure
@@
Figure 8: DCM model forecast (Model 3 simulated time series for combination $n=50$ and ${\mathrm{SN}}=10$: first forty observations are forecasting data). (a) One signal model $g_{1,1,0}$ results. Forecasting data are $t_{i}$, $y_{i}$ and $\sigma_{i}$ ($n=40$ black dots) and forecasted data are ten last observations $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$ ($n=10$ open black dots). Continuous black line of forecasting data model $g(t,\bm{\beta}_{\mathrm{Fore}})$ ends to vertical dotted blue line, where dotted black line of forecasted model $g^{\prime}(t,\bm{\beta}_{\mathrm{Fore}})$ begins. Dotted red line denotes $\pm 3\sigma_{g(t)}$ errors of both models (Equation 30). Residuals of forecasting data model $\epsilon_{i}=y_{i}-g_{i}$ (blue dots), residuals of forecasted data model $\epsilon_{i}=^{\prime}y_{i}^{\prime}-g_{i}^{\prime}$ (open blue dots) and $\pm 3\sigma$ errors of both models (red dotted line) are offset to level -1 (blue dotted line). (b) Two signal model $g_{2,1,0}$ results. Otherwise as in “a”. (c) Three signal model $g_{3,1,0}$ results. Otherwise as in “a”.
3.11 Ill-posed problem
3.11.1 Existence ( $\mathrm{C}_{1}$)
[13] transformed the original function into the frequency domain. The modern DFT time series analysis method transforms the original time series into the frequency domain and gives the best frequency for a pure sine model. [15] presented the LS method, which minimises the differences between the data and the linear model. Our DCM does the same by testing a large number of linear models. The data spacing, even or uneven, is irrelevant for these models. For every chosen DCM model, the total number of tested linear models is
$n_{\mathrm{Lin}}=\binom{n_{\mathrm{L}}}{K_{1}}+(1+n_{\mathrm{B}})\times\binom{n_{\mathrm{S}}}{K_{1}}$, (50)
where the number of tested long and short search frequencies is $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$, respectively. The number of bootstrap samples is $n_{\mathrm{B}}$. In other words, we solve this ill-posed problem by using brute computational force. If the time series contains only zero mean white noise, the Gauß-Markov theorem ensures that a LS fit solution exists for every tested frequency combination.
3.11.2 Uniqueness ( $\mathrm{C}_{2}$)
We make the most of the Gauß-Markov theorem. When the numerical values of the tested frequencies $\bm{\beta}_{I}$ (Equation 9) are fixed, the model becomes linear and the solution for the other free parameters $\bm{\beta}_{II}$ is unique (Equation 2.1). All possible $\bm{\beta}_{I}$ frequency combinations are tested (Equation 12). For every tested frequency combination $\bm{\beta}_{I}$, the linear model gives a unique value for the test statistic $z$ (Equations  16 and 17). From all tested frequency combinations, we select the best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ which minimises $z$. The linear model for this best frequency combination $\bm{\beta}_{\mathrm{I,best}}$ gives unique values for the remaining other free parameters $\bm{\beta}_{\mathrm{II,best}}$. The only goal for the massive DCM search (Equation 50) is to find these unique initial free parameter values $\bm{\beta}_{\mathrm{initial}}=[\bm{\beta}_{\mathrm{I,best}},\bm{\beta}_{\mathrm{II,best}}]$ for the non-linear iteration that gives the unique final free parameter values $\bm{\beta}_{\mathrm{final}}$ (Equation 22).
We use the Fisher-test to compare many different non-linear DCM models against each other. The selection criterion for the best model is unique (Equation 48). The best DCM model is not necessarily the correct model, if this correct model is not among the compared models. The correct model must be able to forecast the future and past data. We formulate the Forecast-test for alternative DCM models (Equation 3.12). The order of Fisher- and Forecast-tests can be reversed. However, the former uses all data, while the latter uses a subset of all data, the forecasting data. In ideal cases, both tests identify the same best and correct model, like in Section 3.12.
3.11.3 Stability ( $\mathrm{C}_{3}$)
There are additional signatures of stability. The $z$ periodogram solution is unique for every tested $\bm{\beta}_{I}$ frequency combination. If these periodograms are continuous and their changes are not irregular (e.g., like in Figs. 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$. Furthermore, the solutions for all seven simulated time series converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7). The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
We conclude that our computational statistical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 3.11.1 - 3.11.3).
3.12 Forecast
@@
There are numerous techniques for forecasting a time series [e.g., 20, 22, 26]. The DCM model $g(t_{i},\bm{\beta})$ can be used to forecast. We divide all data into the forecasting data and the forecasted data. The time points, the observations and the errors of these samples are
  • $n$ forecasting data values $t_{i}$, $y_{i}$ and $\sigma_{i}$
  • $n^{\prime}$ forecasted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$
The DCM gives the best forecasting data model
$g_{i}=g(t_{i},\bm{\beta}_{\mathrm{Fore}}),$ (51)
@@
where $\bm{\beta}_{\mathrm{Fore}}$ are the free parameter values. The $t_{\mathrm{mid,Fore}}\!=\!(t_{n}\!+\!t_{1})/2$ and $\Delta T_{\mathrm{Fore}}\!=\!t_{n}\!-\!t_{1}$ values are computed from the forecasting data time points $t_{i}$. The forecasting model values at any arbitrary time $t$ can be obtained from the $\bm{\beta}_{\mathrm{Fore}}$ and $t_{\mathrm{mid,Fore}}$ $\Delta T_{\mathrm{Fore}}$ values.
The forecasted data model values are
$g^{\prime}_{i}$ $=$ $g(t^{\prime}_{i},\bm{\beta}_{\mathrm{Fore}}),$ (52)
@@
where $t_{\mathrm{mid}}=t_{\mathrm{mid,Fore}}$ and $\Delta T=\Delta T_{\mathrm{Fore}}$. We do not compute “new” $t_{\mathrm{mid}}$ and $\Delta T$ values from the forecasted data time points $t^{\prime}_{i}$ because the correct $g^{\prime}_{i}$ values are obtained only from the $\bm{\beta}_{\mathrm{Fore}}$, $t_{\mathrm{mid,Fore}}$ and $\Delta T_{\mathrm{Fore}}$ combination of the forecasting model (Equation 52). The $n^{\prime}$ forecasted data model residuals
@@
give the forecasted data test statistic
$z_{\mathrm{Fore}}$ $=$ z test statistic for forecasted
data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ (Equation 16 or 17).
@@
This parameter $z_{\mathrm{Fore}}$ measures how well the forecast (Equation 52) obtained from the forecasting data works for the forecasted data. If the DCM detects a new signal from the forecasting data, there are two alternatives:
@@
  • If this new signal is real, the $z_{\mathrm{Fore}}$ value of forecasted data decreases.
@@
  • If this new signal is unreal, the $z_{\mathrm{Fore}}$ value of forecasted data increases.
@@
This “Forecast-test” technique revealed at least five real signals in the sunspot record (Jetsu 2025).
@@
We compute the $z_{\mathrm{Fore}}$ parameter value from the known forecasted data $t^{\prime}_{i}$, $y^{\prime}_{i}$ and $\sigma^{\prime}_{i}$ values (Equation 3.12). Forecasts are possible even if all $t^{\prime}_{i}$, $y^{\prime}_{i}$ or $\sigma^{\prime}_{i}$ forecasted data values are unknown. In this case, the $t_{i}$, $y_{i}$ and $\sigma_{i}$ values of all data can be used as forecasting data. The best DCM model $g(t_{i},\bm{\beta}_{\mathrm{Fore}})$ for all data determines the correct $\bm{\beta}_{\mathrm{Fore}}$, $t_{\mathrm{mid,Fore}}=(t_{n}+t_{1})/2$ and $\Delta T_{\mathrm{Fore}}=t_{n}-t_{1}$ combination of the forecasting model. The $n^{\prime}$ forecasted data time points $t_{i}^{\prime}$ can be created for any arbitrary chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{i}^{\prime}$. The $n^{\prime}$ forecasted $g_{i}^{\prime}=g^{\prime}_{i}(t_{i}^{\prime},\bm{\beta}_{\mathrm{Fore}})$ values are obtained from Equation 52. These $g^{\prime}_{i}$ values can be used, for example, to compute the forecasted mean level
$m_{\mathrm{Fore}}={{1}\over{n^{\prime}}}\sum_{i=1}^{n^{\prime}}g^{\prime}_{i}$ (54)
@@
during the chosen sample window $\Delta T^{\prime}=t_{n}^{\prime}-t_{1}^{\prime}$. Jetsu (2025) used this $m_{\mathrm{Fore}}$ parameter to postdict the known $\Delta T^{\prime}$ time intervals of prolonged solar activity minima, like the Maunder minimum between the years 1640 and 1720. Since the known mean level of all sunspot data was $m$, Jetsu (2025) used these three criteria for correct postdictions of past prolonged activity minima $\Delta T^{\prime}$ time intervals:
@@
  • If DCM detects a real new signal in all data, the $m_{\mathrm{Fore}}$ value decreases.
@@
  • If DCM detects many real signals in all data, the $m_{\mathrm{Fore}}$ value falls below $m$.
@@
  • If DCM detects an unreal new signal in all data, the $m_{\mathrm{Fore}}$ value increases.
@@
We use Model 3 combination $n=50$ and ${\mathrm{SN}}=10$ simulated data (Table 3, column 2) to illustrate the DCM forecasting technique. The black dots in Figures 8a-c are the first $n=40$ forecasting data values $t_{i}$, $y_{i}$ and $\sigma_{i}$. The open black dots denote the $n^{\prime}=10$ forecasted data values $t_{i}^{\prime}$, $y_{i}^{\prime}$ and $\sigma_{i}^{\prime}$. The continuous black line is the forecasting model $g(t)$ and the dotted black line is the forecasted model $g^{\prime}(t)$. The red dotted line shows the $\pm 3\sigma$ errors of both models. The blue dots are the forecasting data $\epsilon_{i}=y_{i}-g_{i}$ residuals and the open blue dots are the forecasted data $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ residuals.
@@
Model 3 is the sum of two $P_{1}=0.16$ and $P_{2}=0.17$ pure sine signals $(K_{1}=2,K_{2}=1)$ superimposed on the constant mean level $M_{0}=1$ $(K_{3}=0)$. We compute the $z$ and $z_{\mathrm{Fore}}$ values for the $g_{1,1,0}$, $g_{2,1,0}$ and $g_{3,1,0}$ models (Table 10). These three models have the same $K_{2}=1$ and $K_{3}=0$ orders as Model 3, but their signal numbers $K_{1}=1,2$ or 3 are different, the $g_{2,1,0}$ model being the correct simulation Model 3.
@@
The correct $g_{2,1,0}$ model gives the smallest $z_{\mathrm{Fore}}=0.88$ value (Table 10). Therefore, it is a better forecasting model than the $g_{1,1,0}$ and $g_{3,1,0}$ models. The one signal $g_{1,1,0}$ model forecast fails because the blue open circles denoting the forecasted data residuals $\epsilon_{i}^{\prime}=y_{i}^{\prime}-g_{i}^{\prime}$ show large deviations from to the blue dotted line offset level of $\epsilon_{i}^{\prime}=-1$ (Figure 8a). The two signal $g_{2,1,0}$ model forecast succeeds because all open blue dots denoting the forecasted data residuals stay close to the blue dotted line offset level $\epsilon_{i}^{\prime}=-1$, as well as inside the red dotted $\pm 3\sigma$ model error limits (Figure 8b).
@@
The three signal model $g_{3,1,0}$ periods are $P_{1}=0.0643\pm 0.0018$, $P_{2}=0.1592\pm 0.0012$ and $P_{3}=0.1716\pm 0.0012$. The amplitudes of these pure sine signals are $A_{1}=0.092\pm 0.025$, $A_{2}=1.879\pm 0.16$ and $A_{3}=1.82\pm 0.15$. The periods and the amplitudes of the two strongest $P_{2}$ and $P_{3}$ signals are correct because they are the same as in Model 3 simulation (Table 3, column 1). Therefore, the $g_{3,1,0}$ forecast appears nearly as good as the $g_{2,1,0}$ forecast because the residuals $\epsilon_{i}^{\prime}$ are close to the offset level of $\epsilon_{i}^{\prime}=-1$ (blue open dots), and these residuals also stay inside the red dotted $\pm 3\sigma$ model error limits (Figure 8c). The amplitude $A_{1}=0.092$ of the third $P_{1}=0.0643$ signal is very low. Due this weak “unreal” $P_{1}$ signal, the $g_{3,1,0}$ model has a larger $z_{\mathrm{Fore}}=0.91$ value than the correct $g_{2,1,0}$ model (Table 10). Finally, we note that the three signal model $g_{3,1,0}$ is not unstable ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}”) although the simulated time series contains only two signals.
For the forecasting data, the extreme Fisher-test critical levels $Q_{\mathrm{F}}<10^{-13}$ confirm that the two signal $g_{2,1,0}$ and the three signal $g_{3,1,0}$ models are certainly better than the one signal $g_{1,1,0}$ model. The forecasting data parameters $n=40$, $\eta_{1}=7$, $\eta_{2}=10$, $\chi_{1}=22.946$ and $\chi_{2}=19.790$ for the simple $g_{2,1,0}$ model and the complex $g_{3,1,0}$ model give the Fisher-test critical level $Q_{\mathrm{F}}=0.28>\gamma_{\mathrm{F}}=0.001$. The $g_{2,1,0}$ model beats the $g_{3,1,0}$ model because the $H_{0}$ hypothesis is not rejected. Hence, the Fisher-test confirms that the $g_{2,1,0}$ model is the best model for the forecasting data. The best DCM model (Fisher-test, $g_{2,1,0}$ ) is also the correct DCM model (Forecast-test, $g_{2,1,0}$). The double-check works for this particular Model 3 simulated time series!
The DFT can not detect the correct frequencies for the forecasting data because the simulated frequencies are “too close” (Equation 37). Therefore, the DFT forecast can not succeed.
@@
The analyses of all seven complex time series indicate that the relative accuracy of amplitude and period estimates are lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 1 - 7). This statistical effect is the same when DCM is applied to the time series of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude forecasts are less accurate than our solar cycle minimum and maximum epoch forecasts (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [25, 17] would explain why correct signal periods have not been detected earlier from the sunspot data.
The DCM analysis results can be double-checked. We use the Fisher-test to identify the best DCM model from all tested DCM models (Equation 48). The correct model may not be among the tested models. Hence, the best model is not necessarily the correct one. This best model must be correct if it passes the Forecast-test (Equation 3.12). This Fisher- and Forecast-test combination double-checks the DCM analysis results. The perfect result is that the best model is the correct model. Therefore, the Forecast-test also prevents overfitting.